How many gallons of milk containing 1% butterfat must be mixed with 30 gallons of milk containing 3% butterfat to obtain a mixture containing 2% butterfat?
Answer by Paul(988)

x+3(30)=2(x+30)
x+90=2x+60
-x=-30
x=30

Hence, about 30 gallons
Paul.